OpenStudy (anonymous):

Find the range of the function for the given domain. f(x)=3x-5;{-3,0,1,5} A.{-14,-5,-2,10} B.{-9,0,3,15} C.{-5,-2,-1,3} D.{0,3,4,8}

OpenStudy (anonymous):

@AnswerMyQuestions @iGreen.

OpenStudy (anonymous):

the range are the y values. each coordinate is (x,y). all you need to do is thist the second number in each coordinate.

OpenStudy (anonymous):

Range is the y-values.

OpenStudy (anonymous):

*list

OpenStudy (anonymous):

The points are in (x, y) So the 2nd number is the y-value. Examples: (3, 6) 6 is the y-value here

OpenStudy (anonymous):

We have to find the whole points.

OpenStudy (anonymous):

i thinking B or A

OpenStudy (anonymous):

plug in -3 for x in the equation and solve for f(x)

OpenStudy (anonymous):

do the same with each x value. each answer you get will be in the range.

OpenStudy (anonymous):

\(f(x)=3x-5\) Plug in -3, 0, 1, and 5: \(f(x)=3(-3)-5\) \(f(x)=-9-5\) \(f(x)=-14\) So our first point is (-3, -14)

OpenStudy (anonymous):

What's the range in (-3, -14)? @idontwantogotosleep

OpenStudy (anonymous):

next just plug in 0. \[f(x)=3(0)-5\]

OpenStudy (anonymous):

No, he just needs to find the range of (-3, -14) and he has his answer.

OpenStudy (anonymous):

would A be correct /.

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

(-3, -14) The range here is -14. And only one option has -14 in it..A.

OpenStudy (anonymous):

oh. :P

OpenStudy (anonymous):

thanks guys , y'all are BIG help :P

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

no problem!